40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 28 20:42:53 2020, Koji, Update, General, Vent Prep 10x
    Reply  Thu Oct 29 12:43:35 2020, gautam, Update, General, Vent Prep preVentAlignment.pngPRMIcarr_PRMalignment.pngALS_ool.pdflowPowerMC.png
       Reply  Thu Oct 29 19:30:56 2020, gautam, Update, General, Vent complete ventValveStatus.pngvent_Oct2020.jpgventValveStatus_overnight.pnglowPowerIMC_afterVent.png
          Reply  Mon Nov 2 12:44:25 2020, gautam, Update, General, ETMY and OMC chamber heavy doors off CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
             Reply  Mon Nov 2 16:46:06 2020, gautam, Update, General, Plan for OMC chamber IMG_2318.JPGIMG_2332.JPG
                Reply  Mon Nov 2 17:13:19 2020, Koji, Update, General, Plan for OMC chamber 
                   Reply  Mon Nov 2 17:32:05 2020, gautam, Update, General, Plan for OMC chamber IMG_2317.JPG
Message ID: 15652     Entry time: Thu Oct 29 19:30:56 2020     In reply to: 15651     Reply to this: 15653
Author: gautam 
Type: Update 
Category: General 
Subject: Vent complete 
  1. Jam nuts - checked that none of the nuts move by hand, which according to Steve, is sufficient. I recorded photos of all of them here.
  2. Lab particle count - at the SP table, I measured 25,910 / cf @ 0.3 um and 1220 /cf  @ 0.5 um. Steve's guidance is that the latter number should be less than 10,000, so no issues there.
  3. Valves - closed all the annuli off, and also closed VM1, VA6 and V1. The procedure calls for VM3 to be opened, presumably so that the RGA remains pumped, but I see no reason why we can't just leave that RGA volume valved off.
  4. Started letting N2 into the main volume at ~3pm local time, by opening VV1. Attachment #1 shows the valve configuration adopted for this vent. Went up to 25 torr and then switched over to the "Ultra Zero" instrument grade air cylinders.
  5. Aimed for 3-5 torr/min rate of pressure increase locally. The full vent trend can be seen in Attachment #2.
  6. Stopped letting air into the main volume when P1a pressure was 700 torr, at which point I disconnected the cylinders from the main volume, and left VV1 open so that the IFO equilibriates to the lab air pressure. I used 4 full cylinders of the instrument grade air, which is par for the course. 
  7. Since we anticipate opening the ETMY and output optics chambers, I also vented those annuli. The final state of the valves as I am leaving it for the night is shown in Attachment #3.
  8. I re-aligned the IMC mirrors so that I could lock the IMC at low power once again. Indeed, IMC trans of ~1400 cts was realized (see Attachment #4), c.f. ~1500 earlier today, I thought this is fine and didn't optimize further. I think the policy is to not lock the IMC in air unless necessary, so I'm leaving the PSL shutter closed for the night.
  9. The ITMs, ETMs and BS were re-aligned such that the Oplev spots are centered. I can see some higher order resonances of the green beams. The AS spot is fringing (the two ND=0.5 filters on the AS CCD camera were removed for better visibility). So this is fine as far as alignment is concerned.

We are now ready to take the doors off. I've already done the basic prep work (loosened bolts, cleaned chamber, carts for tools, fresh ameristat on portable HEPAs etc).

Quote:

If everything else looks good, I'll start letting the dry N2 into the main volume after lunch.

Attachment 1: ventValveStatus.png  41 kB  Uploaded Thu Oct 29 20:38:50 2020  | Hide | Hide all | Show all
ventValveStatus.png
Attachment 2: vent_Oct2020.jpg  216 kB  Uploaded Thu Oct 29 22:50:12 2020  | Show | Hide all | Show all
Attachment 3: ventValveStatus_overnight.png  41 kB  Uploaded Thu Oct 29 22:51:59 2020  | Hide | Hide all | Show all
ventValveStatus_overnight.png
Attachment 4: lowPowerIMC_afterVent.png  54 kB  Uploaded Thu Oct 29 22:52:04 2020  | Hide | Hide all | Show all
lowPowerIMC_afterVent.png
ELOG V3.1.3-